### v1

@@ Abstract
Differentially private distributed mean estimation (DP-DME) is a fundamental building block in privacy-preserving federated learning, where a central server estimates the mean of $d$-dimensional vectors held by $n$ users while ensuring $(\epsilon,\delta)$-DP. Local differential privacy (LDP) and distributed DP with secure aggregation (SecAgg) are the most common notions of DP used in DP-DME settings with an untrusted server. LDP provides strong resilience to dropouts, colluding users, and malicious server attacks, but suffers from poor utility. In contrast, SecAgg-based DP-DME achieves an $O(n)$ utility gain over LDP in DME, but requires increased communication and computation overheads and complex multi-round protocols to handle dropouts and malicious attacks. In this work, we propose CorDP-DME, a novel DP-DME mechanism that spans the gap between DME with LDP and distributed DP, offering a favorable balance between utility and resilience to dropout and collusion. CorDP-DME is based on correlated Gaussian noise, ensuring DP without the perfect conditional privacy guarantees of SecAgg-based approaches. We provide an information-theoretic analysis of CorDP-DME, and derive theoretical guarantees for utility under any given privacy parameters and dropout/colluding user thresholds. Our results demonstrate that (anti) correlated Gaussian DP mechanisms can significantly improve utility in mean estimation tasks compared to LDP – even in adversarial settings – while maintaining better resilience to dropouts and attacks compared to distributed DP.
[…]
1 Introduction
@@
Distributed mean estimation (DME) is a fundamental building block in a number of applications ranging from federated learning [1, 2], distributed stochastic gradient descent [3, 4, 5, 6] to distributed sensor network computations [7]. Differentially private DME (DP-DME) refers to the setting where a central server aims to estimate the mean of $d$ dimensional vectors held by $n$ distributed users while ensuring differential privacy (DP) [8, 9] of the users’ vectors. The utility of DP-DME is measured by the mean squared error (MSE) between the estimate at the server and the true mean. DP-DME has been studied under multiple notions of DP: (i) Central DP (CDP) [10], where a trusted server collects user vectors, computes the mean, and adds noise to ensure DP; (ii) Local DP (LDP), where each user independently perturbs their vector before sending to a potentially malicious server, and (iii) Distributed DP [11], where users perturb vectors locally and use secure aggregation (SecAgg) [12, 13] to ensure the server only views the sum of perturbed vectors [14, 15, 16, 17].
LDP and distributed DP using SecAgg are the most common approaches for DME that avoid the requirement of a trusted server – see Table 1 for a comparison. DME with LDP requires nearly no coordination among users and, consequently, maintains privacy even if an arbitrary number of users collude with the server. Additionally, it provides a graceful degradation of utility with user dropouts – that is, the MSE increases gradually as the number of users who drop out increases. However, the privacy resilience of LDP-based DME comes at a large utility cost: relative to CDP, LDP-based DME increases MSE by a factor of $O(n)$ [18].
Distributed DP via SecAgg, on the other hand, requires perfect coordination across users and a complex multi-party cryptographic protocol performed over multiple rounds. This approach has two key features: (i) it ensures that the server learns only the sum of the user’s vectors (and nothing else) and (ii) it allows for an MSE that is achieved by CDP, and hence an $O(n)$ factor reduction in the MSE as compared to LDP. In effect, because of (i), each user only needs to add enough noise to ensure privacy of the summed vectors, which enables the server’s input – as well as the resulting MSE – to be statistically identical to the CDP case. However, the improvement in MSE comes at a coordination cost. SecAgg requires perfectly (anti) correlated random noise from a finite field to mask the users’ vectors to ensure that the server can only access the sum of the users’ vectors. This makes the method less resilient to dropouts in terms of utility and vulnerable to malicious server attacks that involve fake client registrations and strategic connection drops [19]. Notably, the protocol can be made resilient to dropouts and collusions up to fixed thresholds on the number of dropouts and colluding users. However, unlike LDP, this resilience requires multiple rounds of communication and increased shared randomness among users. Moreover, when the collusion and dropout thresholds are exceeded, DP-DME with SecAgg results in sudden drops in privacy and utility, respectively.
In this work, we develop novel DP-DME mechanisms that simultaneously retain the key advantages of both LDP and SecAgg. Specifically, we target techniques that: (i) like LDP, can be performed with one round without requiring complex cryptographic approaches to handle dropouts and collusions, (ii) like SecAgg, provide MSE comparable to CDP, (iii) are resilient like LDP, in that they yield graceful degradation of utility with increasing number of collusions and dropouts irrespective of the thresholds, (iv) maintains the required level of privacy up to the colluding user threshold as in both LDP and distributed DP with SecAgg, and graceful decay of privacy beyond the threshold. State-of-the-art distributed DP methods for DME use cryptographic protocols such as SecAgg that operate over a finite field and ensure perfect privacy of the inputs and perfect reconstruction of the sum of the inputs. However, these methods ultimately provide DP guarantees to users’ real-valued vectors and approximate reconstructions of the mean. We introduce an alternative DP-DME setting that is end-to-end differentially private.
Our starting point (see Sec. 3) is a novel interpretation of one of the core ideas of SecAgg: utilizing (anti) correlated noise to reduce the overall MSE without compromising on privacy. Unlike SecAgg, which requires correlated noise over a finite field, we consider adding correlated Gaussian noise directly to real-valued vectors held by users.
@@
Figure 1: MSE with different numbers of responding users for $n=100$, $\epsilon=2$, $\delta=10^{-5}$. CorDP-DME spans the gap between DME with LDP and CDP.
@@
We couple this starting point with a DP framework for quantifying the impact of arbitrary noise correlations on privacy, particularly in face of user dropouts and collusion. We apply this framework to develop a simple, one-round DP-DME protocol called CorDP-DME. After an initial offline phase to establish shared randomness, CorDP-DME uses an optimized correlated Gaussian mechanism among users and provides the utility of CDP without resorting to SecAgg (See Table 1). In the presence of dropouts and collusions, CorDP-DME achieves significantly reduced MSE relative to LDP yet retains its flexibility, resilience, and simplicity as compared to SecAgg. Notably, CorDP-DME spans the gap between the two extremes of LDP and distributed DP with SecAgg (Fig. 1).
1.1 Our Contributions
@@
  • We introduce a DP-DME model that allows users to execute privacy mechanisms that are arbitrarily correlated with each other. The proposed model assumes a malicious server that attempts to compromise users’ privacy, and considers dropouts and colluding users. The generalized notion of DP for DME introduced in this work covers both LDP and distributed DP with SecAgg as special cases (Section 2).
@@
  • We provide a novel geometric interpretation of the use of (anti) correlated privacy mechanisms in DP-DME, which demonstrates how correlated mechanisms can be used to achieve utility gains over independent noise mechanisms used by LDP (Section 3).
@@
  • We perform an information-theoretic analysis on the correlated Gaussian mechanism for DP-DME to determine the optimum noise parameters and the decoder at the server that minimizes the MSE for any given privacy parameters and dropout/colluding user thresholds (Section 5).
@@
  • We propose CorDP-DME, a single-round DP-DME mechanism that spans the gap between DME with LDP and distributed DP with SecAgg in terms of privacy-utility trade-offs and resilience to dropouts and attacks (Sections 4 and 5).
LDP [20] Distributed DP w/ SecAgg [12, 16] CorDP-DME
Communication Single round Multiple rounds Single round

### v2

@@ Abstract
Differentially private distributed mean estimation (DP-DME) is a fundamental building block in privacy-preserving federated learning, where a central server estimates the mean of $d$-dimensional vectors held by $n$ users while ensuring $(\epsilon,\delta)$-DP. Local differential privacy (LDP) and distributed DP with secure aggregation (SA) are the most common notions of DP used in DP-DME settings with an untrusted server. LDP provides strong resilience to dropouts, colluding users, and adversarial attacks, but suffers from poor utility. In contrast, SA-based DP-DME achieves an $O(n)$ utility gain over LDP in DME, but requires increased communication and computation overheads and complex multi-round protocols to handle dropouts and attacks. In this work, we present a generalized framework for DP-DME, that captures LDP and SA-based mechanisms as extreme cases. Our framework provides a foundation for developing and analyzing a variety of DP-DME protocols that leverage correlated privacy mechanisms across users. To this end, we propose CorDP-DME, a novel DP-DME mechanism based on the correlated Gaussian mechanism, that spans the gap between DME with LDP and distributed DP. We prove that CorDP-DME offers a favorable balance between utility and resilience to dropout and collusion. We provide an information-theoretic analysis of CorDP-DME, and derive theoretical guarantees for utility under any given privacy parameters and dropout/colluding user thresholds. Our results demonstrate that (anti) correlated Gaussian DP mechanisms can significantly improve utility in mean estimation tasks compared to LDP – even in adversarial settings – while maintaining better resilience to dropouts and attacks compared to distributed DP.
Index Terms:
Differential privacy, distributed mean estimation, noise correlation
I Introduction
@@
Distributed mean estimation (DME) is a fundamental building block in a number of applications ranging from federated learning [1, 2], distributed stochastic gradient descent [3, 4, 5, 6] to distributed sensor network computations [7]. Differentially private DME (DP-DME) refers to the setting where a central server aims to estimate the mean of $d$ dimensional vectors held by $n$ distributed users while ensuring differential privacy (DP) [8, 9] of the users’ vectors. The utility of DP-DME is measured by the mean squared error (MSE) between the estimate at the server and the true mean. DP-DME has been studied under multiple notions of DP:
  1. (i)
    Central DP (CDP) [10], where a trusted server collects user vectors, computes the mean, and adds noise to ensure DP;
  2. (ii)
    Local DP (LDP) [11, 12], where each user independently perturbs their vector before sending to a potentially adversarial server that performs aggregation;
  3. (iii)
    Distributed DP with secure aggregation (SA) [13, 14, 15, 16, 17], where users perturb vectors locally and use a SA protocol [18, 19, 20] to ensure the server only views the sum of the perturbed vectors, and
  4. (iv)
    Shuffle DP [21, 22, 23], where the locally perturbed vectors of the users are shuffled by a trusted shuffler before sending to the server for aggregation.
LDP and distributed DP with SA are the most common approaches to DP-DME that eliminate the need for a trusted third party. [^1] Other DP-DME approaches exist, employing distributed DP in decentralized (graph-based) settings [24, 25]. A comparison of our work with these approaches is provided in Section I-B. LDP and SA offer distinct advantages for DME, such as reduced overhead in LDP and enhanced utility in SA. In this work, we address a fundamental question:
How should we design a DP-DME mechanism that achieves a higher utility than LDP while significantly reducing the overhead relative to SA?
In order to answer this question, we first provide more details on the advantages and limitations of each method.
LDP-based DME protocols [11, 26] are simple, single-round methods that require nearly no coordination among users. It assumes the strongest threat model in DP-DME with an adversarial server and no constraints on the number of users colluding with the server or dropouts. However, the robust privacy guarantees of LDP-based DME come at a significant utility cost. For the same level of privacy, the MSE resulted by LDP-based DME is higher than the MSE of CDP by a factor of $O(n)$ [27].
In contrast, distributed DP via SA achieves the same privacy-utility trade-off as CDP-based DME by relying on perfect coordination and synchronization among users, along with a complex multi-party cryptographic protocol executed over multiple rounds. The key feature of SA is that the server only learns the sum of the users’ uploads. In other words, SA ensures that the data received by the server is independent of the user’s inputs, conditioned on the sum. This allows users to apply less noise to their private vectors than LDP, thus improving utility.
The utility gain in distributed DP with SA comes at a high coordination cost. Practical implementations of SA (e.g., the SecAgg [^2] A detailed description of the SecAgg protocol is provided in Appendix H, along with a simple four-user example illustrating the key ideas. protocol in [18]) operate in two main phases: an offline phase for pre-processing (shared randomness generation) and an online phase where user-vectors are uploaded and aggregated by the server. SecAgg relies on perfectly (anti) correlated pair-wise random noise, uniformly sampled from a finite field for each pair of users. For this, SecAgg employs the Diffie-Hellman key exchange among all participating users during the offline phase. The use of finite field noise requires a costly multi-round dropout recovery mechanism to handle even a single dropout in the online phase. To facilitate this, each user distributes components of all pairwise random seeds as secret shares with every other user during the offline phase. Such complex dropout recovery mechanisms result in increased communication, computational overhead, synchronization, and the need for extensive shared randomness among users. As discussed in [16], the increased cost of multi-round interactions between users and the server can thwart the efficient training of large machine learning models, particularly in federated learning settings with unreliable communication links. Furthermore, SecAgg enforces strict constraints on the maximum number of dropouts and colluding users, and fails if these limits are exceeded. SecAgg’s multiple communication rounds between the server and users also make the protocol vulnerable to malicious server attacks [28].
In this paper, we develop a DP-DME framework that addresses the shortcomings of both LDP-based and SA-based DME. In particular, we identify the following underlying factors contributing to inefficiencies.
  • Increased MSE with LDP: LDP-based mechanisms are independent across users, leading to increased MSE.
  • Complexity and overheads of SA: The complexity and overheads of SA protocols largely stems from its multi-round dropout handling mechanism,
  • Strict dropout thresholds in SA: The lack of flexibility in SA-based DP-DME, due to strict limits on dropouts and colluding users, as well as the reliance on complex multi-round dropout recovery mechanisms, arise from the use of random noise from a finite field.
Our key insight is that, ultimately, the privacy guarantee of SA-based DP-DME is $(\epsilon,\delta)$-DP, which obviates overly conservative cryptographic measures such as sampling random noise from a finite field. Building on this observation, we define a DP-DME framework that is end-to-end DP and simplifies dropout handling, resulting in protocols that are more flexible and resilient to attacks and dropouts, compared to DP-DME with SA. Specifically, we target techniques that: (i) like LDP, can be performed with one round without requiring complex cryptographic approaches to handle dropouts and collusions, (ii) like SA, provide MSE comparable to CDP, (iii) like LDP, are resilient to privacy attacks, and do not cease to operate beyond dropout/colluding user thresholds (see Table I [^3] Table I provides a comparison of the two main approaches to DP-DME and CorDP-DME. For SA-based DP-DME, SecAgg [18] is selected as the underlying SA protocol for fair comparison based on the common assumptions made. Further comparisons between CorDP-DME and other SA protocols are discussed in Section I-B3.).
@@
Figure 1: MSE of LDP, CDP and CorDP-DME with different numbers of responding users for a setting with $n=100$, $\epsilon=2$, $\delta=10^{-5}$. CorDP-DME coincides with CDP when all users respond. All three mechanisms coincide when only one user responds. CorDP-DME always outperforms LDP. In general, CorDP-DME spans the gap between DME with LDP and CDP.
@@
Our starting point is a novel geometric interpretation (see Sec. III) of the privacy-utility trade-off in DP-DME, which reflects how (anti) correlated mechanisms can outperform independent (LDP) mechanisms. Our analysis demonstrates that SA, in essence, represents an extreme case within the broader spectrum of DP-DME mechanisms described by this geometric framework. Unlike SA, which requires correlated noise over a finite field, we consider adding correlated Gaussian noise directly to real-valued vectors held by users. The use of correlated Gaussian noise leads to single-round DP-DME protocols [^4] Single-round protocols refer to those with only one round in the (online) DME phase, excluding the (offline) noise generation phase. that are robust to dropouts and colluding users. We couple this starting point with a DP framework for quantifying the impact of arbitrary noise correlations on privacy, particularly in the face of user dropouts and collusion. We apply this framework to develop a simple, one-round DP-DME protocol called CorDP-DME. After an initial offline phase to establish shared randomness, CorDP-DME uses an optimized correlated Gaussian mechanism among users and provides the utility of CDP without resorting to SA in absence of dropouts (See Table I). In the presence of dropouts and collusions, CorDP-DME achieves significantly reduced MSE relative to LDP yet retains its flexibility, resilience, and simplicity compared to SA. Notably, CorDP-DME spans the gap between the two extremes of LDP and distributed DP with SA (Fig. 1).
From a broader theoretical perspective, we present a generalized framework for DP-DME that serves as a foundation for developing a range of DP-DME protocols with varying properties. LDP-based and SA-based DME represent extreme points within this spectrum, each with distinct characteristics. Our framework enables the design of protocols that combine the key advantages of both extremes – such as enhanced utility and reduced complexity and overheads – while also supporting theoretical insights into fundamental limits and trade-offs.
I-A Our Contributions
@@
  • We introduce a generalized distributed DP framework for DME that accounts for users that execute privacy mechanisms that are arbitrarily correlated with each other. The proposed model considers both dropouts and colluding users. DME with LDP and distributed DP with SA are extreme cases of the proposed generalized DP-DME framework (Section II).
@@
  • We provide a novel geometric interpretation of the privacy-utility trade-off in DP-DME, which demonstrates how (anti) correlated privacy mechanisms can outperform independent privacy mechanisms used by LDP (Section III).
@@
  • We perform an information-theoretic analysis on the correlated Gaussian mechanism for DP-DME. We derive the optimum noise parameters and the decoding strategy at the server that minimizes the MSE for a target $(\epsilon,\delta)$ privacy parameters and number of dropout/colluding user thresholds. We also provide a converse result that establishes the optimum noise covariance structure that minimizes the MSE of unbiased mean estimates (Section IV-A).
@@
  • Building on the information-theoretic analysis, we propose CorDP-DME, a single-round DP-DME mechanism that spans the gap between DME with LDP and distributed DP with SA in terms of privacy-utility trade-offs and resilience to dropouts and attacks (Section IV).
LDP [12] Distributed DP w/ SecAgg [18, 14] CorDP-DME (ours)
Rounds in protocol Single round Multiple rounds Single round
MSE: dropouts $\leq\ p$ $O\left(\frac{d}{(n-p)\min\{\epsilon,\epsilon^{2}\}}\right)$ $O\left(\frac{d}{(n-p)^{2}\min\{\epsilon,\epsilon^{2}\}}\right)$ $O\left(\frac{dp}{n(n-p)\min\{\epsilon,\epsilon^{2}\}}\right)$
Computation User: $O(d)$ Server: $O(d)$ User: $O(n^{2}+dn)$ Server: $O(dn^{2})$ User: $O(dn)$ Server: $O(d)$
Communication: pre-processing None required User: $O(n)$ key exchanges $O(n)$ secret shares Server: $O(n^{2})$ key exchanges $O(n^{2})$ secret shares User: $O(n)$ key exchanges Server: $O(n^{2})$ key exchanges
Communication: DME phase User: $O(d)$ Server: $O(dn)$ User: $O(d+n)$ Server: $O(dn+n^{2})$ User: $O(d)$ Server: $O(dn)$
Storage User: $O(d)$ Server: $O(d)$ User: $O(d+n)$ Server: $O(d+n^{2})$ User: $O(d+n)$ Server: $O(d)$